JEF meeting, January 17, 2025

From GlueXWiki
Jump to: navigation, search

Connection

Meeting time and location:

Agenda

Preparation of parts for ECAL -- Sasha

-Updated design of the LMS light source with 25 LEDS. A new PCB with 25 LEDs is ready. Prepare mounting flange to the integrating sphere (machine shop), 3D print housing (Bobby)
-More powerful fans will be installed in the central region (Tim)

-Installation of the FCAL2 detector & integration to GlueX -- Sasha, All

  • Implementing ECAL into the GlueX DAQ [1]
- rocecal3 crate was fixed [2] (initial DAQ implementation [3])
- implementing flash ADC scalers [4][5]
- translation table for the ECAL (finalize/check translation DB, generate xml file, import to the CCDB) - done [6]
- implementing and verifying ECAL LMS trigger [7]
- implementing and verifying the ECAL physics trigger. First event reconstructed using the GlueX software [8]
- Start implementing VTP trigger modules. Requires modification of firmware. The trigger crates may be unavailable for 1 - 1.5 months.
- The VTP was installed in the D2-1-TOP crate, IP address assigned, and can be booted now
- New version of DAQ, VTP included as a DAQ component [9]
  • Installation of scintillating paddles for cosmic tests [10] [11][12]
  • ECAL HV interlock check [13]
  • Cable map of the ECAL LG modules after installing ECAL [14]
  • Reinstalling cables on the ECAL LV distribution board [15]
  • Safety briefings - Vlad, Sasha
  • Detector installation page
  • Commissioning north part of the ECAL (instrumented with 2 temporary HV distribution boards) [16][17][18][19]
  • Commissioning of the south bottom region [20]
  • Commissioning of the south top region [21]
  • Commissioning the entire detector using the LMS (signal waveforms) [22][23]
  • Fix LV issues (north top region) [24]
  • First cosmic event in the ECAL [25]
  • Temperature scan in the ECAL divider region [26][27][28][29]
Temperature of the NPS detector
2 temperature sensors in the divider region connected, fan included into the detector interlock [30] -- Josh, Bobby
Meeting with Tim: More powerful fans will be installed in the inner region. Beam pipe can be cooled, need specs and technical details (decided to put on hold). Can operate the chillers at smaller temperatures.
  • Repairing a few ECAL LMS optical fibers [31][32]
  • Preparing for the installation of the insulation on the detector face
- Removing Tedlar from the face of the detector [33]
- ECAL flash ADC scalers before removing detector face cover [34]
  • Preparing for the run -- Sasha
Page for shift Takers
ECAL in the Experimental Safety Assessment Document (ESAD)
EPAS for Maintenance of the Hall D ECAL (Jlab-PTW-5236)

Restoring PS test setup [35]

ECAL commissioning -- Sasha, All

Laveen update[[36]]

Calibrations

-First look on cosmic data -- Zhikun
-Adjusting HVs
-pi0 calib -- ig

Analysis/Simulation

Minutes

Attendees: Liping, Bill, Sasha, Igal, Zhikun, Vlad, Olga, Drew, Phil, S.T.

  • Chris S has assembled new PCB with 25 LEDs for LMS
    • more powerful driver, brighter LEDs
    • Needs to be mounted to integrating sphere
  • 3D parts for reference PMT assembly in the works
  • Fans in inner region of insert will be replaced
  • Measured LMS waveforms for all channels in insert
    • Found some low amplitude/distorted waveforms
    • Loose connectors fixed
    • One channel had baseline problems -> flash ADC replaced
    • One channel had amplitude changing over time -> divider replaced
    • All channels now operational
  • Zhikun looked at cosmic data, fitted peak positions with Landau distributions
    • Set all peak positions to 8 counts -> gain equalization -> new HV settings
  • Can now control HV using CSS
  • Fan now in HV interlock
  • 4 temperature sensors working, also in HV interlock
  • Temperature in insert region/on the face of the crystals still a concern
    • Reducing HV from fixed value of 1.04 kV to the channel-dependent values determined by the cosmic ray procedure reduced temperature on the face when HV is on from 30C to 26C
    • Relaxation time/ramp up time when HV turned off/on is on the order of several hours
      • Measure resolution right after turning on and then later after stabilization to study impact?
      • Study shower shape with 5x5 prototype in PS arm?
    • Gradient along z through the modules observed
      • Light output sensitive to temperature -> impact on effective shower shape
    • Considering cooling beam pipe: more material in beam hole region
    • Considering lowering chiller temperature
    • NPS temperature measurements give some indication of what we might expect
  • Igal presented calibration procedures for ECAL using π0s
    • Assumes initial gain calibration using cosmics/Compton events and timing calibration
      • Trigger needs to have gains equalized in the hardware at ~10% level
    • Tested with simulated 2π0 events
    • Boundary between ECAL and lead glass needs to be treated carefully due to shower leakage
  • Calibration using Compton events
    • Turn on all tagger counters? For low energy
    • Two arm or one arm? One arm would be easier, but what about the background?